# Emergent Dominance Hierarchies in Reinforcement Learning Agents

### v6

@@ 6.2. Comparison to animal behavior
Inspired by the geometric study of intransitive policies in Czarnecki et al. (2020), we visually compare the occurrence of intransitive components (cycles) between Chicken Coop populations and mice populations, shown in Figure 6. In Williamson et al. (2017), 20 captive populations of 12 male CD-1 mice each were observed for aggressive behavior for 22 days. [^3] The data was conveniently aggregated by Strauss et al. (2022). We compare these results to those from $L=30$ populations of $N=12$ Chicken Coop agents. In order to find the maximum similarity between the natural and artificial populations, we use a learning rate of $3\times 10^{-5}$. We make this comparison by using a metric we define here as rank linearity [^4] The metric of rank linearity is similar to a metric called dominance certainty, introduced in McCowan et al. (2022) and measured in rhesus macaque breeding groups., which expresses how likely each dominance rank is to be occupied by a single agent, rather than a $K$-way tie. For a single population, we can describe rank linearity as a boolean; for example, in the third population from the top in Figure 4, ranks $\{0,4,5\}$ are linear, while ranks $\{1,2,3\}$ are non-linear as they are part of a 3-way tie between agents $1$, $3$ and $5$. For a set of populations, we define rank linearity as a number in $[0,1]$ which is the probability of a rank being linear in a sample population.
@@ 6.4. Transmitting dominance hierarchies to new populations
In previous subsections we regarded Chicken Coop agents as the focal objects of study, while we considered the dominance hierarchies that they form as a pattern in their behavior. In this subsection we propose flipping this perspective on its head: We consider dominance hierarchies as the focal object while the agents are used as vessels that communicate the dominance hierarchy that they learned to a new population, promoting a cultural evolution paradigm in artificial intelligence (Team et al., 2022; Tomlinson, 2009).
@@ 7. Discussion and future work
We thank the following colleagues for their advice and support in writing this paper: David Aha, Ivan Chase, Edgar Duéñez-Guzmán, Errol King, Joel Leibo, Olof Leimar, David Manheim, Markov, Georg Ostrovski, Jérémy Perret, Saul Pwanson, Venkatesh Rao and Eli Strauss.

### v7

@@ 6.2. Comparison to animal behavior
Inspired by the geometric study of intransitive policies in Czarnecki et al. (2020), we visually compare the occurrence of intransitive components (cycles) between Chicken Coop populations and mice populations, shown in Figure 6. In Williamson et al. (2017), 20 captive populations of 12 male CD-1 mice each were observed for aggressive behavior for 22 days. [^3] The data was conveniently aggregated by Strauss et al. (2022). We compare these results to those from $L=30$ populations of $N=12$ Chicken Coop agents. In order to find the maximum similarity between the natural and artificial populations, we use a learning rate of $3\times 10^{-5}$. We make this comparison by using a metric we define here as rank linearity [^4] The metric of rank linearity is similar to a metric called dominance certainty, introduced in McCowan et al. (2022) and measured in rhesus macaque breeding groups., which expresses how likely each dominance rank is to be occupied by a single agent, rather than a $K$-way tie. For a single population, we can describe rank linearity as a boolean; for example, in the third population from the top in Figure 4, the top rank (rank 0) and the two bottom ranks (ranks 4 and 5) are occupied by agents 4, 0 and 2 respectively. Therefore, we consider ranks 0, 4 and 5 as linear. The remaining ranks are 1, 2 and 3, and they cannot be directly correlated with any agent, because agents 1, 3 and 5 are in a three-way tie. Therefore, we consider ranks 1, 2 and 3 as non-linear. For sets of populations, we define rank linearity as a number in $[0,1]$ which is the probability of a rank being linear in a sample population.
@@ 6.4. Transmitting dominance hierarchies to new populations
In previous subsections we regarded Chicken Coop agents as the focal objects of study, while we considered the dominance hierarchies that they form as a pattern in their behavior. In this subsection we propose flipping this perspective on its head: We consider dominance hierarchies as the focal object while the agents are used as vessels that communicate the dominance hierarchy that they learned to a new population, promoting a cultural evolution paradigm in artificial intelligence (Team, 2022; Tomlinson, 2009).
@@ 7. Discussion and future work
We thank our colleagues for their advice and support in writing this paper: David Aha, Ivan Chase, Edgar Duéñez-Guzmán, Errol King, Joel Leibo, Olof Leimar, David Manheim, Markov, Georg Ostrovski, Jérémy Perret, Saul Pwanson, Venkatesh Rao and Eli Strauss.